# ul. Pasteura 5, 02-093 Warsaw, Poland

### v1

@@ Abstract
The Stochastic Gravitational Wave Background (SGWB) from cosmic superstrings offers one of the few known possibilities to test String Theory within current experimental reach. However, in order to be compatible with the existing constraints, the tension of a cosmic superstring network is required to lie several orders of magnitude below the Planck scale. This is naturally realized in string compactifications where the volume of the extra dimensions is parametrically large (in string units). We estimate the GW spectrum arising from cosmic superstrings in such scenarios, providing analytical formulae as well as numerical results. Crucially, we do so within a fully-fledged string cosmology, taking into account various modified cosmological epochs (such as kination or early matter domination) induced by the presence of moduli and a time-dependent string tension. We show that part of the spectrum generically lies within reach of LISA and ET, with a large class of models predicting a characteristic drop in the amplitude which may be robustly probed by LISA. The corresponding signal would encode information on the dynamics of moduli and reheating. On the other hand, the ultra-high frequency part of the spectrum can be significantly enhanced by a long, early phase of kination with time-varying tension, yielding a spectral index unique to this set-up.
@@ 1 Introduction
The observation of gravitational waves (GWs) has started a new era in observational cosmology LIGOScientific:2016aoc . So far, most efforts have concentrated on the detection of events that are astrophysical in nature, such as black holes or neutron star mergers Bailes:2021tot . It is however possible that the discovery of (putative) stochastic GW backgrounds of cosmological origin might ignite a similar revolution in fundamental physics. Within this context, cosmic strings are a particularly interesting target, being one of the very few universal predictions [^1] Their existence as objects in the spectrum of the theory is guaranteed. Whether they are produced in the early Universe is model-dependent. of String Theory which might be accessible to current experimental methods. The eventual detection of the GW background from a network of cosmic (super)strings could give a spectacular confirmation of String Theory, which is very likely beyond the reach of planned collider experiments. In order to be consistent with observations, the tension $\mu$ of superstrings in our universe would have to be parametrically smaller than the Planck scale, $\mu\lesssim 10^{-11}-10^{-12}\,M_{P}^{2}$, where the strongest bounds come from GW searches with Pulsar Timing Arrays (PTAs) NANOGrav:2023gor ; EPTA:2023fyk ; Reardon:2023gzh ; Xu:2023wog at nHz frequencies and throughout the paper we use the dimensionally reduced Planck mass $M_{P}=1/\sqrt{8\pi G}\simeq 2.4\times 10^{18}$ GeV. [^2] Large scale interferometers such as LVK also give significant constraints, although these may be relaxed by early-time modifications to standard cosmology Ellis:2023tsl . This will indeed be the case in the scenarios we study.
@@
In String Theory, a very natural and model-independent way to realize this is in large volume compactifications, where the size of the extra dimensions is very large in units of $M_{P}$ Balasubramanian:2005zx ; Conlon:2005ki . These scenarios and their phenomenology have been studied intensely, as the large volumes also realize the limit of control where computations can be reliably performed within the supergravity approximation. It is interesting to notice that realistic values for the extra-dimensional volume in such constructions fall exactly in the right ballpark to reproduce the recently observed signal from PTAs NANOGrav:2023gor ; EPTA:2023fyk ; Reardon:2023gzh ; Xu:2023wog in terms of GWs from cosmic superstrings. [^3] See Conzinu:2024cwl for another possible explanation of the PTA signal with a string cosmology origin, in the Pre-Big-Bang scenario Gasperini:1992em . While an astrophysical explanation of the signal (super massive black hole binaries) is the most conservative NANOGrav:2023gor ; Ellis:2023dgf ; EPTA:2023xxk , [^4] See NANOGrav:2023hvm ; Figueroa:2023zhu ; Ellis:2023oxs ; Ratzinger:2020koh for an in-depth comparison between many cosmological possibilities. cosmic superstrings also provide a good fit to the data Ellis:2023tsl ; Avgoustidis:2025svu . This is to be contrasted with the situation for gauge strings, which cannot provide a good fit to the amplitude and the slope of the signal NANOGrav:2023hvm ; Ellis:2023tsl ; Ellis:2020ena ; Blasi:2020mfx ; Blanco-Pillado:2021ygr .  [^5] In particular, the predicted slope is not steep enough to accommodate the data.
@@
In addition to revealing the existence of (super)strings, the detection of any cosmic string GW background (i.e. not limited to string theory ones) could at the same time serve as a powerful probe of early universe cosmology. Indeed, one particularly appealing feature of cosmic strings is that they evolve in a self-similar way (dubbed as the scaling regime Albrecht:1984xv ; Bennett:1987vf ; Allen:1990tv ) throughout cosmic history, thus emitting GWs throughout their whole evolution, unlike other cosmological sources [^6] An obvious exception being the stochastic GW background from inflation. (such as phase transitions) whose emission is localized in time. Therefore, their signal would be spread across the whole observable spectrum and potentially visible to experiments of very different types, ranging from PTAs to large scale interferometers such as LIGO-Virgo-KAGRA (LVK) KAGRA:2013rdx , Einstein Telescope Sathyaprakash:2012jk (ET) and LISA Audley:2017drz ; Blanco-Pillado:2024aca .
@@
It is well known (see e.g. Gouttenoire:2019kij and references therein) that the spectral index of the GW background depends on the equation of state of the Universe at the time that the GWs are sourced. Thus, if cosmic strings were present in our Universe, these could well be studied through measurement of stochastic backgrounds of gravitational waves in different ranges of frequencies. More concretely, the GW spectrum sourced by a network of cosmic strings during radiation domination is flat. Deviations from radiation domination thus induce boosts or dips in the spectrum (see e.g. (Gouttenoire:2019kij, )). Departures from the “standard cosmology” between inflation and radiation domination have been recently studied in the context of large-volume string compactifications Conlon:2022pnx ; Apers:2022cyl ; Revello:2023hro ; Apers:2024ffe ; Conlon:2024hgw ; Conlon:2024ene (see Cicoli:2023opf for a review), where the physics of moduli seem to suggest that the early Universe had a long, interesting history with epochs with various equations of state.
@@
An advantage unique to GWs is that they are also able to probe epochs before Big Bang Nucleosynthesis, so far unconstrained by observation Allahverdi:2020bys ; Cicoli:2024bwq and where stringy effects are expected to be even more relevant Apers:2024ffe . In trying to access information about such early times, however, an obvious difficulty is posed by the fact the GW signals sourced very early on will peak at very high frequencies. This is an unavoidable consequence of causality; the wavelength of a GW must lie within the Hubble radius at the time it is emitted, resulting in a lower bound on the frequency. Interesting frequency values for early universe cosmology roughly span the $10\,{\rm{kHz}}$$-100\,{\rm{GHz}}$ band, in what is known as the Ultra High Frequency (UHF) GW regime. As a limitation/opportunity, these are currently beyond the range of current detectors, which cannot cover frequencies much larger than 10 kHz. However, no astrophysical sources are known to emit GWs in this region, making it a particularly compelling target: any measured signal would have to be of cosmological origin or new physics in the astrophysical regime, such as primordial black holes or exotic objects. UHF gravitational waves are a very active topic of research, and extensive theoretical and experimental efforts are being spent to investigate new detection techniques (See Aggarwal:2020olq ; Roshan:2024qnv ; Aggarwal:2025noe for reviews).
@@
Finally, let us also make a general point which does not hinge on the specific scenarios discussed in our paper. One of the most generic consequences of String Theory (aside from the existence of strings) is the fact that there exist no free parameters, and all couplings arise from expectation values of scalar fields. It follows that the properties of cosmic strings, whether field theoretic or fundamental, can also depend on the moduli in a stringy setting. To make a simple example, the fundamental string tension $m_{s}^{2}$ in a generic compactification will depend on the stabilised values of the volume and the dilaton, as reviewed below. In scenarios where the moduli are evolving on cosmological timescales, such as those mentioned in the first paragraph, this will invariably lead to cosmic string networks with a varying tension. For this reason, a main goal of the paper will be to perform a model-indepedent analysis of how the usual properties of cosmic string networks are modified in the case of a time-dependent tension, and how it affects their GW emission. We will present explicit examples, both a toy version and a more realistic scenario, where this gives rise to a spectral index that cannot be accommodated by strings with constant tension.
@@
The paper is structured as follows. In Section 2, we review cosmic strings in String Theory, mostly focusing on the properties of fundamental string networks and the effect of a varying tension. Section 3 provides a brief introduction to large volume compactifications and the typical cosmological histories arising in that setting. In particular, we discuss how moduli dynamics induce periods of time-varying tension and matter domination, where the latter gives rise to a prediction for the LISA band in a large class of models. The computation of GW spectrum is contained in Section 4 where, under the assumption of a scaling regime (shown to occur with time-varying tension in Revello:2024gwa ), we provide novel analytic tools for the computation of GW spectra with constant and time-dependent tensions, accommodating the different dependences on the intercommutation probability proposed in the literature. Section 5 is devoted to a few selected benchmark scenarios and their observational consequences, most notably including a concrete prediction for a drop of the spectrum in the LISA band. Finally, we conclude in Section 6 with some future prospects.
@@ 2 Cosmic superstrings
Cosmic strings encompass a variety of one-dimensional, topological defects that may have formed in the early universe. In Quantum Field Theory (QFT), they are realized as string-like, solitonic field configurations with a finite energy per unit length, given by the tension $\mu$ [^7] We are neglecting possible IR divergences, which may be regulated on physical grounds. Perhaps the simplest example is that of a quartic scalar theory with a spontaneously broken $\rm{U}(1)$ symmetry, where the winding around a string in spacetime corresponds to winding around the valley of the potential. In String Theory, they can rather arise as microscopic, fundamental strings that have been stretched by Hubble expansion, for example during inflation. Of course, the two pictures are compatible, as field theory strings can naturally appear in low-energy limits of string theory corresponding to a QFT. In that case, the field-theoretic singularity at the core of the string is UV completed into either a fundamental string or a wrapped brane Reece:2023czb .
@@
2.1 Microscopic properties
Stable strings in the early universe can form macroscopic networks, which are the source of the gravitational waves considered in this paper. Such networks can consist of fundamental (F-) strings, D1 branes and bound states of the two, known as $(p,q)$-strings. [^8] A $(p,q)$ string is precisely a bound state of $p$ F-strings and $q$ D-strings. However, in the region of frequencies we are interested in, the largest contribution to the GW signal comes from the F-strings only (See Avgoustidis:2025svu for the most up-to-date analysis of a multi-component $(p,q)$-string network). Therefore, we will assume only F-strings are present. The two fundamental, microscopic parameters determining the evolution of a fundamental cosmic string network are the string tension $\mu$ and the intercommutation probability $P$. The string tension is determined by the string scale in the case of fundamental strings, or by the scalar vev of some symmetry breaking field in the scalar case. The intercommutation probability $P$, on the other hand, is the probability for two long string segments to form a loop when they collide. In field theory $P\simeq 1$, while in a string theoretic setting it usually suppressed by the string coupling $g_{s}$. This should already make it clear that cosmic superstrings networks will behave rather differently compared to their field-theoretic counterpart, with a crucial impact for observations mainly due to their modified evolution. Interestingly, it is precisely thanks to these differences that the cosmic superstring interpretation of the recently measured signal from PTAs NANOGrav:2023gor ; EPTA:2023fyk ; Reardon:2023gzh ; Xu:2023wog is favored with respect to that of (stable) field theory cosmic strings Ellis:2023tsl .
2.1.1 String tension
A distinguishing feature of the String Theory case is that the tension of the strings is not necessarily constant, as it can inherit a dependence on the various moduli of a given compactification Revello:2024gwa . This is a specific instance of a more general principle, according to which energy scales in string theory are not fixed a priori, but rather determined by dynamical processes. As an example, the fundamental string tension in a type II compactification (without warping) is given by
$\mu=\frac{1}{2\pi\alpha^{\prime}}=m_{s}^{2}=\frac{g_{s}^{1/2}}{2\mathcal{V}}M_{P}^{2}$ (1)
where $\sqrt{\alpha^{\prime}}$ is the fundamental unit of length in string theory (the Regge slope), $m_{s}$ the string scale, $g_{s}$ the string coupling [^9] In terms of the string dilaton $\varphi$, $g_{s}=e^{\varphi}$. and $\mathcal{V}$ the volume of the extra dimensions (in the Einstein frame and units of $2\pi\sqrt{\alpha^{\prime}}$). In the above, the string scale $m_{s}$ is related to $M_{P}$ by the expectation values of the dilaton and volume modulus respectively. If such fields exhibit a non-trivial time dependence during the evolution of the universe, the tension of the strings will vary with time, translating to a significant impact on the evolution of the string network Revello:2024gwa . Moreover, this is not unique to String Theory strings - the same phenomenon is expected to be generic for field theory strings within a stringy UV completion. Indeed, time-varying moduli can also induce time variations of scalar field vevs (rather than fundamental scales like $m_{s}$), and thus of QFT string tensions. One of the questions explored in this work is whether such a distinctive phenomenology may leave imprints on the stochastic gravitational wave background.
From an observational perspective, the tension $\mu$ is constrained to be orders of magnitude below $M_{P}^{2}$ from a variety of sources, implying a parametric suppression of the string scale. The analysis of CMB anisotropies gives the robust limit $G\mu<10^{-7}-10^{-8}$ Charnock:2016nzm , while Pulsar Timing Array (PTA) searches for stochastic GW backgrounds NANOGrav:2023hvm ; EPTA:2023fyk ; Reardon:2023gzh ; Xu:2023wog ; NANOGrav:2023gor ; EPTA:2023sfo ; EPTA:2023fyk can improve the bound by many orders of magnitude, $G\mu\lesssim 10^{-11}-10^{-12}$ Blanco-Pillado:2021ygr ; Ellis:2023tsl ; Marfatia:2023fvh ; Avgoustidis:2025svu . This already rules out cosmic strings in heterotic string compactifications Witten:1985fp , where the tension is essentially fixed in terms of the GUT fine structure constant as $G\mu\sim\frac{\alpha_{\rm{GUT}}}{16\pi}\sim 10^{-3}$. On the other hand, (1) shows how low tensions are not difficult to obtain in the type II setting. In particular, it is clear that a parametrically small string tension can most naturally be achieved in the weakly coupled limit of large compactification volumes (and eventually small string couplings). For this reason, it is very natural to focus on scenarios with large extra dimensions, where $\mathcal{V}\gg 1$.
Large Extra Dimensions were initially suggested as a purely phenomenological way to ameliorate or solve the hierarchy problem Arkani-Hamed:1998jmv , but later found a string theoretic incarnation in terms of the Large Volume Scenario (LVS) Balasubramanian:2005zx ; Conlon:2005ki  [^10] It must be remarked, however, that the similarity is only qualitative. While in both cases the extra-dimensional radius is parametrically larger than the string length, in the LVS construction it is well below the $R\sim 1\,\rm{mm}$ advocated in Arkani-Hamed:1998jmv . and its generalizations. The LVS is a construction arising from flux compactifications of type IIB, and is one of the better established scenarios of moduli stabilization, featuring in particular a mechanism to stabilize the volume modulus at exponentially large values. Moreover, it is very appealing from the point of view of phenomenology as the exponentially large volume easily allows to engineer hierarchies between different scales in Nature - such as the one between $m_{s}$ and $M_{P}$. Realistic values for the stabilized volume and dilaton in LVS are in the range $\mathcal{V}\sim 10^{6}-10^{13}$ and $g_{s}\sim 0.1$, resulting a string tension $G\mu\sim 10^{-9}-10^{-16}$. These values not only allow to evade the constraints mentioned above, but are in the right ballpark to be tested by upcoming GW detectors.
For these reasons, we shall take the LVS and related constructions as an interesting benchmark in the study of gravitational waves from string theory. Another reason, to be explored in detail in Section 3, is that scenarios with steep runaway potentials towards large volume strongly suggest modifications to the $\Lambda$CDM models of cosmology in the very early universe which can further modify the spectrum of GW from cosmic strings. All in all, this paints the compelling and consistent picture of a well-motivated corner of String Theory which allows for observationally viable (but not yet excluded) cosmic superstrings and at the same time can exhibit smoking-gun signatures of its string theoretic origin, in the form of a time-dependent string tension.
@@ 2.1.1 String tension
For completeness, let us also mention another possible mechanism to obtain a low, fundamental string tension. This can happen if the string scale is locally suppressed by a significant amount of warping, which can happen if the compactification space (usually taken to be a Calabi-Yau manifold) develops throats. In that case, the metric can locally be described by Dasgupta:1999ss ; Giddings:2001yu ; Giddings:2005ff :
$ds^{2}=\left(1+\frac{e^{4\mathcal{A}(y)}}{\mathcal{V}^{2/3}}\right)^{-1/2}g_{\mu\nu}dx^{\mu}dx^{\nu}+\left(1+\frac{e^{4\mathcal{A}(y)}}{\mathcal{V}^{2/3}}\right)^{1/2}g_{mn}dy^{n}dy^{m},$ (2)
@@
$\mu=\frac{1}{2\pi\alpha^{\prime}}\left(1+\frac{e^{4\mathcal{A}(y_{0})}}{\mathcal{V}^{2/3}}\right)^{-1/2}\simeq\frac{g_{s}^{1/2}}{2\mathcal{V}^{2/3}}e^{-2\mathcal{A}(y_{0})}M_{P}^{2}\,,$ (3)
with a clear exponential suppression. [^11] This expression corrects a factor of $2\pi$ in Revello:2024gwa and uses the 10D Einstein frame expression for the warp factor Giddings:2005ff . While this mechanism is relatively common in string compactifications (for example in the KKLT construction Kachru:2003aw ), many of the details are model-dependent, and we shall not pursue this case further.

### v2

# Cosmic superstrings in large volume compactifications: PTAs, LISA and time-varying tension

@@ Abstract
The Stochastic Gravitational Wave Background (SGWB) from cosmic superstrings offers one of the few known possibilities to test String Theory within current experimental reach. However, in order to be compatible with the existing constraints, the tension of a cosmic superstring network is required to lie several orders of magnitude below the Planck scale. This is naturally realized in string compactifications where the volume of the extra dimensions is parametrically large (in string units). We estimate the GW spectrum arising from cosmic superstrings in such scenarios, providing analytical formulae as well as numerical results. Crucially, we do so within a semi-realistic string cosmology scenario, taking into account various modified cosmological epochs (such as kination or early matter domination) induced by the presence of moduli and a time-dependent string tension. We show that part of the spectrum generically lies within reach of LISA and ET, with a large class of models predicting a characteristic drop in the amplitude which may be robustly probed by LISA. The corresponding signal would encode information on the dynamics of moduli and reheating. On the other hand, the ultra-high frequency part of the spectrum can be significantly enhanced by a long, early phase of kination with time-varying tension, yielding a spectral index unique to this set-up.
@@ 1 Introduction
The observation of gravitational waves (GWs) has started a new era in observational cosmology [1]. So far, most efforts have concentrated on the detection of events that are astrophysical in nature, such as black holes or neutron star mergers [2]. It is however possible that the discovery of (putative) stochastic GW backgrounds of cosmological origin might ignite a similar revolution in fundamental physics. Within this context, cosmic strings are a particularly interesting target, being one of the very few universal predictions [^1] Their existence as objects in the spectrum of the theory is guaranteed. Whether they are produced in the early Universe is model-dependent. of String Theory which might be accessible to current experimental methods. The eventual detection of the GW background from a network of cosmic (super)strings could give a spectacular confirmation of String Theory, which is very likely beyond the reach of planned collider experiments. In order to be consistent with observations, the tension $\mu$ of superstrings in our universe would have to be parametrically smaller than the Planck scale, $\mu\lesssim 10^{-11}-10^{-12}\,M_{P}^{2}$, where the strongest bounds come from GW searches with Pulsar Timing Arrays (PTAs) [3, 4, 5, 6] at nHz frequencies and throughout the paper we use the reduced Planck mass $M_{P}=1/\sqrt{8\pi G}\simeq 2.4\times 10^{18}$ GeV. [^2] Large scale interferometers such as LVK also give significant constraints, although these may be relaxed by early-time modifications to standard cosmology [7, 8]. This will indeed be the case in the scenarios we study.
@@
In String Theory, a very natural and model-independent way to realize this is in large volume compactifications, where the size of the extra dimensions is very large in units of $M_{P}$ [9, 10]. These scenarios and their phenomenology have been studied intensely, as the large volumes also realize the limit of control where computations can be reliably performed within the supergravity approximation. It is interesting to notice that realistic values for the extra-dimensional volume in such constructions fall exactly in the right ballpark to reproduce the recently observed signal from PTAs [3, 4, 5, 6] in terms of GWs from cosmic superstrings. [^3] See [11] for another possible explanation of the PTA signal with a string cosmology origin, in the Pre-Big-Bang scenario [12]. While an astrophysical explanation of the signal (super massive black hole binaries) is the most conservative [3, 13, 14], [^4] See [15, 16, 17, 18] for an in-depth comparison between many cosmological possibilities. cosmic superstrings also provide a good fit to the data [7, 19]. This is to be contrasted with the situation for gauge strings, which cannot provide a good fit to the amplitude and the slope of the signal [15, 7, 20, 21, 22].  [^5] In particular, the predicted slope is not steep enough to accommodate the data.
@@
In addition to revealing the existence of (super)strings, the detection of any cosmic string GW background (i.e. not limited to string theory ones) could at the same time serve as a powerful probe of early universe cosmology. Indeed, one particularly appealing feature of cosmic strings is that they evolve in a self-similar way (dubbed as the scaling regime [23, 24, 25]) throughout cosmic history, thus emitting GWs throughout their whole evolution, unlike other cosmological sources [^6] An obvious exception being the stochastic GW background from inflation. (such as phase transitions) whose emission is localized in time. Therefore, their signal would be spread across the whole observable spectrum and potentially visible to experiments of very different types, ranging from PTAs to large scale interferometers such as LIGO-Virgo-KAGRA (LVK) [26], Einstein Telescope [27] (ET) and LISA [28, 29].
@@
It is well known (see e.g. [30] and references therein) that the spectral index of the GW background depends on the equation of state of the Universe at the time that the GWs are sourced. Thus, if cosmic strings were present in our Universe, these could well be studied through measurement of stochastic backgrounds of gravitational waves in different ranges of frequencies. More concretely, the GW spectrum sourced by a network of cosmic strings during radiation domination is flat. Deviations from radiation domination thus induce boosts or dips in the spectrum (see e.g. [30]). Departures from the “standard cosmology” between inflation and radiation domination have been recently studied in the context of large-volume string compactifications [31, 32, 33, 34, 35, 36] (see [37] for a review), where the physics of moduli seem to suggest that the early Universe had a long, interesting history with epochs with various equations of state.
@@
An advantage unique to GWs is that they are also able to probe epochs before Big Bang Nucleosynthesis, so far unconstrained by observation [38, 39] and where stringy effects are expected to be even more relevant [34]. In trying to access information about such early times, however, an obvious difficulty is posed by the fact the GW signals sourced very early on will peak at very high frequencies. This is an unavoidable consequence of causality; the wavelength of a GW must lie within the Hubble radius at the time it is emitted, resulting in a lower bound on the frequency. Interesting frequency values for early universe cosmology roughly span the $10\,{\rm{kHz}}$$100\,{\rm{GHz}}$ band, in what is known as the Ultra High Frequency (UHF) GW regime. As a limitation/opportunity, these are currently beyond the range of current detectors, which cannot cover frequencies much larger than 10 kHz. However, no astrophysical sources are known to emit GWs in this region, making it a particularly compelling target: any measured signal would have to be of cosmological origin or new physics in the astrophysical regime, such as primordial black holes or exotic objects. UHF gravitational waves are a very active topic of research, and extensive theoretical and experimental efforts are being spent to investigate new detection techniques (See [40, 41, 42] for reviews).
@@
Finally, let us also make a general point which does not hinge on the specific scenarios discussed in our paper. One of the most generic consequences of String Theory (aside from the existence of strings) is the fact that there exist no free parameters, and all couplings arise from expectation values of scalar fields. It follows that the properties of cosmic strings, whether field theoretic or fundamental, can also depend on the moduli in a stringy setting. To make a simple example, the fundamental string tension $m_{s}^{2}$ in a generic compactification will depend on the stabilised values of the volume and the dilaton, as reviewed below. In scenarios where the moduli are evolving on cosmological timescales, such as those mentioned in the first paragraph, this will invariably lead to cosmic string networks with a varying tension. For this reason, a main goal of the paper will be to perform a model-independent analysis of how the usual properties of cosmic string networks are modified in the case of a time-dependent tension, and how it affects their GW emission. We will present explicit examples, both a toy version and a more realistic scenario, where this gives rise to a spectral index that cannot be accommodated by strings with constant tension.
@@
The paper is structured as follows. In Section 2, we review cosmic strings in String Theory, mostly focusing on the properties of fundamental string networks and the effect of a varying tension. Section 3 provides a brief introduction to large volume compactifications and the typical cosmological histories arising in that setting. In particular, we discuss how moduli dynamics induce periods of time-varying tension and matter domination, where the latter gives rise to a prediction for the LISA band in a large class of models. The computation of GW spectrum is contained in Section 4 where, under the assumption of a scaling regime (shown to occur with time-varying tension in [43]), we provide novel analytic tools for the computation of GW spectra with constant and time-dependent tensions, accommodating the different dependences on the intercommutation probability proposed in the literature. Section 5 is devoted to a few selected benchmark scenarios and their observational consequences, most notably including a concrete prediction for a drop of the spectrum in the LISA band. Finally, we conclude in Section 6 with some future prospects.
@@ 2 Cosmic superstrings
Cosmic strings encompass a variety of one-dimensional, topological defects that may have formed in the early universe. In Quantum Field Theory (QFT), they are realized as string-like, solitonic field configurations with a finite energy per unit length, given by the tension $\mu$. [^7] We are neglecting possible IR divergences, which may be regulated on physical grounds. Perhaps the simplest example is that of a quartic scalar theory with a spontaneously broken $\rm{U}(1)$ symmetry, where the winding around a string in spacetime corresponds to winding around the valley of the potential. In String Theory, they can rather arise as microscopic, fundamental strings that have been stretched by Hubble expansion, for example during inflation. Of course, the two pictures are compatible, as field theory strings can naturally appear in low-energy limits of string theory corresponding to a QFT. In that case, the field-theoretic singularity at the core of the string is UV completed into either a fundamental string or a wrapped brane [44].
The main focus of this work will be on cosmic superstrings, which can consist of either fundamental (F-)strings, or Dpp branes wrapped around a (p1)(p-1)- dimensional cycle. For simplicity, we will only consider the case of fundamental F-strings, as they give the dominant contribution to the GW spectrum in the frequency range we are interested in. In this section we intend to give a brief overview of their phenomenological properties, relevant to the computation of gravitational waves. The interested reader may also wish to consult Section 2 of [43], as well as some of the original references on the topic [45, 46, 47, 48, 49, 50].
2.1 Microscopic properties
Stable strings in the early universe can form macroscopic networks, which are the source of the gravitational waves considered in this paper. Such networks can consist of fundamental (F-) strings, D1 branes and bound states of the two, known as $(p,q)$-strings. [^8] A $(p,q)$ string is precisely a bound state of $p$ F-strings and $q$ D-strings. However, in the region of frequencies we are interested in, the largest contribution to the GW signal comes from the F-strings only (See [19] for the most up-to-date analysis of a multi-component $(p,q)$-string network). Therefore, we will assume only F-strings are present. The two fundamental, microscopic parameters determining the evolution of a fundamental cosmic string network are the string tension $\mu$ and the intercommutation probability $P$. The string tension is determined by the string scale in the case of fundamental strings, or by the scalar vev of some symmetry breaking field in the scalar case. The intercommutation probability $P$, on the other hand, is the probability for two long string segments to form a loop when they collide. In field theory $P\simeq 1$, while in a string theoretic setting it usually suppressed by the string coupling $g_{s}$. This should already make it clear that cosmic superstrings networks will behave rather differently compared to their field-theoretic counterpart, with a crucial impact for observations mainly due to their modified evolution. Interestingly, it is precisely thanks to these differences that the cosmic superstring interpretation of the recently measured signal from PTAs [3, 4, 5, 6] is favored with respect to that of (stable) field theory cosmic strings [7].
2.1.1 String tension
A distinguishing feature of the String Theory case is that the tension of the strings is not necessarily constant, as it can inherit a dependence on the various moduli of a given compactification [43]. This is a specific instance of a more general principle, according to which energy scales in string theory are not fixed a priori, but rather determined by dynamical processes. As an example, the fundamental string tension in a type II compactification (without warping) is given by
$\mu=\frac{1}{2\pi\alpha^{\prime}}=m_{s}^{2}=\frac{g_{s}^{1/2}}{2\mathcal{V}}M_{P}^{2}$ (2.1)
where $\sqrt{\alpha^{\prime}}$ is the fundamental unit of length in string theory (the Regge slope), $m_{s}$ the string scale, $g_{s}$ the string coupling [^9] In terms of the string dilaton $\varphi$, $g_{s}=e^{\varphi}$. and $\mathcal{V}$ the volume of the extra dimensions (in the Einstein frame and units of $2\pi\sqrt{\alpha^{\prime}}$). In the above, the string scale $m_{s}$ is related to $M_{P}$ by the expectation values of the dilaton and volume modulus respectively. If such fields exhibit a non-trivial time dependence during the evolution of the universe, the tension of the strings will vary with time, translating to a significant impact on the evolution of the string network [43]. Moreover, this is not unique to String Theory strings - the same phenomenon is expected to be generic for field theory strings within a stringy UV completion. Indeed, time-varying moduli can also induce time variations of scalar field vevs (rather than fundamental scales like $m_{s}$), and thus of QFT string tensions. One of the questions explored in this work is whether such a distinctive phenomenology may leave imprints on the stochastic gravitational wave background.
From an observational perspective, the tension $\mu$ is constrained to be orders of magnitude below $M_{P}^{2}$ from a variety of sources, implying a parametric suppression of the string scale. The analysis of CMB anisotropies gives the robust limit $G\mu<10^{-7}-10^{-8}$ [51], while Pulsar Timing Array (PTA) searches for stochastic GW backgrounds [15, 4, 5, 6, 3, 52, 4] can improve the bound by many orders of magnitude, $G\mu\lesssim 10^{-11}-10^{-12}$ [22, 7, 53, 19]. This already rules out cosmic strings in heterotic string compactifications [45], where the tension is essentially fixed in terms of the GUT fine structure constant as $G\mu\sim\frac{\alpha_{\rm{GUT}}}{16\pi}\sim 10^{-3}$. On the other hand, (2.1) shows how low tensions are not difficult to obtain in the type II setting. In particular, it is clear that a parametrically small string tension can most naturally be achieved in the weakly coupled limit of large compactification volumes (and eventually small string couplings). For this reason, it is very natural to focus on scenarios with large extra dimensions, where $\mathcal{V}\gg 1$.
Large Extra Dimensions were initially suggested as a purely phenomenological way to ameliorate or solve the hierarchy problem [54, 55], but later found a string theoretic incarnation in terms of the Large Volume Scenario (LVS) [9, 10] [^10] It must be remarked, however, that the similarity is only qualitative. While in both cases the extra-dimensional radius is parametrically larger than the string length, in the LVS construction it is well below the $R\sim 1\,\rm{mm}$ advocated in [54]. and its generalizations. The LVS is a construction arising from flux compactifications of type IIB, and is one of the better established scenarios of moduli stabilization, featuring in particular a mechanism to stabilize the volume modulus at exponentially large values. Moreover, it is very appealing from the point of view of phenomenology as the exponentially large volume easily allows to engineer hierarchies between different scales in Nature - such as the one between $m_{s}$ and $M_{P}$. Realistic values for the stabilized volume and dilaton in LVS are in the range $\mathcal{V}\sim 10^{6}-10^{13}$ and $g_{s}\sim 0.1$, resulting a string tension $G\mu\sim 10^{-9}-10^{-16}$. These values not only allow to evade the constraints mentioned above, but are in the right ballpark to be tested by upcoming GW detectors.
@@ 2.1.1 String tension
For completeness, let us also mention another possible mechanism to obtain a low, fundamental string tension. This can happen if the string scale is locally suppressed by a significant amount of warping, which can happen if the compactification space (usually taken to be a Calabi-Yau manifold) develops throats. In that case, the metric can locally be described by [56, 57, 58]:
$ds^{2}=\left(1+\frac{e^{4\mathcal{A}(y)}}{\mathcal{V}^{2/3}}\right)^{-1/2}g_{\mu\nu}dx^{\mu}dx^{\nu}+\left(1+\frac{e^{4\mathcal{A}(y)}}{\mathcal{V}^{2/3}}\right)^{1/2}g_{mn}dy^{n}dy^{m},$ (2.2)
@@
$\mu=\frac{1}{2\pi\alpha^{\prime}}\left(1+\frac{e^{4\mathcal{A}(y_{0})}}{\mathcal{V}^{2/3}}\right)^{-1/2}\simeq\frac{g_{s}^{1/2}}{2\mathcal{V}^{2/3}}e^{-2\mathcal{A}(y_{0})}M_{P}^{2}\,$, (2.3)
with a clear exponential suppression. [^11] This expression corrects a factor of $2\pi$ in [43] and uses the 10D Einstein frame expression for the warp factor [58]. While this mechanism is relatively common in string compactifications (for example in the KKLT construction [59]), many of the details are model-dependent, and we shall not pursue this case further.